\begin{thebibliography}{1}

\bibitem{Adam2014}
 D.P. Kingma, and J. Ba. "Adam: A method for stochastic optimization." arXiv preprint arXiv:1412.6980 (2014)

\bibitem{garcia2018}
  G. Garcia-Hernando, S. Yuan, S. Baek, and T.-K. Kim. First- Person Hand Action Benchmark with RGB-D Videos and 3D Hand Pose Annotations. In CVPR, 2018

\bibitem{sahbiICPR2020} H. Sahbi. "Learning Connectivity with Graph Convolutional Networks." 25th ICPR. IEEE, 2021.

\bibitem{refref10} C. Feichtenhofer, A. P., and A. Zisserman. Convolutional Two-Stream Network Fusion for Video Action Recognition. CVPR, pages 1933-1941, 2016. 8

\bibitem{refref39}  E.Ohn-Barand, M.M.Trivedi. Hand Gesture Recognition in Real Time for Automotive Interfaces: A Multimodal Vision- Based Approach and Evaluations. IEEE TITS, 15(6):2368--2377, 2014.


  
\bibitem{sahbi00009} M. Jiu and H. Sahbi. "Laplacian deep kernel learning for image annotation." 2016 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP). IEEE, 2016.



  
   \bibitem{refref40} 
  O. Oreifej and Z. Liu. HON4D: Histogram of Oriented 4D Normals for Activity Recognition from Depth Sequences. In CVPR, pages 716-723, June 2013.
   \bibitem{refref41}  
  H. Rahmani and A. Mian. 3D Action Recognition from Novel Viewpoints. In CVPR, pages 1506--1515, June 2016.

  \bibitem{Zhua2016} 
  W. Zhu, C. Lan, J. Xing, W. Zeng, Y. Li, L. Shen, and X. Xie. Co-occurrence feature learning for skeleton based action recognition using regularized deep LSTM networks In AAAI, volume 2, page 6, 2016.

\bibitem{Sahbi2011} H.~Sahbi, J.-Y. Audibert, and R.~Keriven, ``Context-dependent kernels for  object classification,'' \emph{IEEE Transactions on Pattern Analysis and   Machine Intelligence}, vol.~33, pp. 699--708, 2011.

  
  \bibitem{refref59} 
  M. Zanfir, M. Leordeanu, and C. Sminchisescu. The Moving Pose: An Efficient 3D Kinematics Descriptor for Low-Latency Action Recognition and Detection. In ICCV, pages 2752--2759, 2013.

  \bibitem{Vemulapalli2014}
  R. Vemulapalli, F. Arrate, and R. Chellappa. Human action recognition by representing 3D skeletons as points in a Lie group. In CVPR 2014.

\bibitem{ref27} N. Bourdis, D. Marraud, and H. Sahbi,  Spatio-temporal interaction for aerial video change detection, in IGARSS, 2012, pp. 2253--2256


  
  \bibitem{Du2015}  
  Y. Du, W. Wang, and L. Wang. Hierarchical recurrent neural network for skeleton based action recognition. In CVPR 2015.

  \bibitem{refref61}  
  X. Zhang, Y. Wang, M. Gou, M. Sznaier, and O. Camps. Efficient Temporal Sequence Comparison and Classification Using Gram Matrix Embeddings on a Riemannian Manifold. In CVPR, pages 4498--4507, 2016
  
 \bibitem{refref11}  
   G. Garcia-Hernando and T.-K. Kim. Transition Forests: Learning Discriminative Temporal Transitions for Action Recognition. In CVPR, pages 407--415, 2017.

\bibitem{Jiu2017}   M.~Jiu and H.~Sahbi, ``Nonlinear deep kernel learning for image annotation,'' \emph{IEEE  Transactions on Image Processing}, vol. 26(4), 2017.

   

  \bibitem{refref18}   
  J. Hu, W. Zheng, J. Lai, and J. Zhang. Jointly Learning Heterogeneous Features for RGB-D Activity Recognition. In CVPR 2015

  \bibitem{Huangcc2017} 
  Z. Huang and L. V. Gool. A Riemannian Network for SPD Matrix Learning. In AAAI, pages 2036–2042, 2017


\bibitem{ref00007777} F.  Fleuret and H.  Sahbi. "Scale-invariance of support vector machines based on the triangular kernel." 3rd International Workshop on Statistical and Computational Theories of Vision. 2003.


  
  \bibitem{ref23} 
  Z. Huang, J. Wu, and L. V. Gool. Building Deep Networks on Grassmann Manifolds. In AAAI, pages 3279–3286, 2018

\bibitem{Krizhevsky2012}
  A.~Krizhevsky, I.~Sutskever, and G.~E. Hinton, \newblock ``Imagenet classification with deep convolutional neural networks,'' \newblock in {NIPS} 2012.

\bibitem{He2016}
K.~He, X.~Zhang, S.~Ren, and J.~Sun,
\newblock ``Deep residual learning for image recognition,''
\newblock in {\em CVPR}, 2016, pp. 770--778.



\bibitem{sahbi00001}  H. Sahbi, "Lightweight Connectivity In Graph Convolutional Networks For Skeleton-Based Recognition." 2021 IEEE International Conference on Image Processing (ICIP). IEEE, 2021.


\bibitem{HuangCVPR2017}
G.~Huang, Z.~Liu, L.~v.~d. Maaten, and K.~Q. Weinberger,
\newblock ``Densely connected convolutional networks,''
\newblock in {\em CVPR}, 2017, pp. 2261--2269.

\bibitem{He2017}
He, Kaiming, et al. "Mask r-cnn." Proceedings of ICCV, 2017.

\bibitem{zhang2020}
  Zhang, Ziwei, Peng Cui, and Wenwu Zhu. "Deep learning on graphs: A survey." IEEE TKDE (2020).


\bibitem{ref00005} N. Boujemaa, F. Fleuret, V. Gouet, H. Sahbi (2004, January). Visual content extraction for automatic semantic annotation of video news. In the proceedings of the SPIE Conference, San Jose, CA (Vol. 6).

  

\bibitem{Ronneberger2015}
O. Ronneberger, P. Fischer, and T. Brox. "U-net: Convolutional networks for biomedical image segmentation." International Conference on Medical image computing and computer-assisted intervention. Springer, 2015.


\bibitem{Bruna2013}
J. Bruna, W. Zaremba, A. Szlam, Y. LeCun. Spectral networks and locally connected networks on graphs. arXiv:1312.6203 (2013)




\bibitem{ref00007} H. Sahbi, L. Ballan, G. Serra, A. DelBimbo (2012). Context-dependent logo matching and recognition. IEEE Transactions on Image Processing, 22(3), 1018-1031.


\bibitem{Henaff2015}
M. Henaff, J. Bruna, Y. LeCun. Deep convolutional networks on graph structured data. arXiv preprint arXiv:1506.05163 (2015)



\bibitem{kipf17} 
TN. Kipf, M. Welling. Semi-supervised classification with graph convolutional networks. In ICLR, 2017

\bibitem{Levie2018}
R. Levie, F. Monti, X. Bresson, M.M. Bronstein. Cayleynets: Graph
convolutional neural networks with complex rational spectral filters. IEEE
Transactions on Signal Processing 67(1), 97--109 (2018)

\bibitem{Sahbi20088888} M. Ferecatu and H. Sahbi. "TELECOMParisTech at ImageClefphoto 2008: Bi-Modal Text and Image Retrieval with Diversity Enhancement." CLEF (Working Notes). 2008.


\bibitem{Li2018}
R. Li, S. Wang, F. Zhu, J. Huang. Adaptive graph convolutional neural
networks. In AAAI, 2018.

\bibitem{Bresson16}
 M. Defferrard, X. Bresson, P. Vandergheynst. Convolutional Neural Networks on graphs with Fast Localized Spectral Filtering. In NIPS, 2016


\bibitem{sahbiiccv17}
  H. Sahbi. Coarse-to-fine deep kernel networks. Proceedings of the IEEE International Conference on Computer Vision, 1131-1139, 2017.


 
\bibitem{Gori2005}
M. Gori, G. Monfardini, F. Scarselli. A new model for learning in
graph domains. In IEEE IJCNN, vol. 2, pp. 729--734, 2005.

\bibitem{Micheli2009}
A. Micheli. Neural network for graphs: A contextual constructive
approach. IEEE TNN 20(3), 498-511 (2009)

\bibitem{Wu2019}
Z. Wu, S. Pan, F. Chen, G. Long, C. Zhang, P.S. Yu. A comprehensive
survey on graph neural networks. arXiv:1901.00596 (2019).


\bibitem{sahbikpca06} H. Sahbi.  Kernel PCA for similarity invariant shape recognition. Neurocomputing 70 (16-18), 3034-3045


\bibitem{Hamilton2017}
W. Hamilton, Z. Ying, J. Leskovec. Inductive representation learning
on large graphs. In NIPS. pp. 1024--1034 (2017)

\bibitem{spectral1997}
Chung, Fan RK, and Fan Chung Graham. Spectral graph theory. No. 92. American Mathematical Soc., 1997.



\bibitem{sahbi00005} H. Sahbi. "Kernel-based Graph Convolutional Networks." 2020 25th International Conference on Pattern Recognition (ICPR). IEEE, 2021.




\bibitem{attention2019}
Knyazev, Boris, Graham W. Taylor, and Mohamed Amer. "Understanding attention and generalization in graph neural networks." Advances in neural information processing systems 32 (2019).
\bibitem{dropconnect2013}
Wan, Li, et al. "Regularization of neural networks using dropconnect." International conference on machine learning. PMLR, 2013.

\bibitem{thiemert2005applying} S.~Thiemert, H.~Sahbi, and M.~Steinebach, ``Applying interest operators in  semi-fragile video watermarking,'' in \emph{Security, Steganography, and   Watermarking of Multimedia Contents VII}, vol. 5681.\hskip 1em plus 0.5em   minus 0.4em\relax International Society for Optics and Photonics, 2005, pp.  353--363.

  
\bibitem{DBLP:conf/cvpr/HuangLMW18}
Gao Huang, Shichen Liu, Laurens van~der Maaten, and Kilian~Q. Weinberger,
\newblock ``Condensenet: An efficient densenet using learned group
  convolutions,''
\newblock in {\em {CVPR}}, 2018.

\bibitem{DBLP:conf/cvpr/SandlerHZZC18}
M.~Sandler, A.~G. Howard, M.~Zhu, A.~Zhmoginov, and L.{-}C. Chen,
\newblock ``Mobilenetv2: Inverted residuals and linear bottlenecks,''
\newblock in {\em {CVPR}}, 2018.



\bibitem{DBLP:journals/corr/HowardZCKWWAA17}
A.~G. Howard, M.~Zhu, B.~Chen, D.~Kalenichenko, W.~Wang, T.~Weyand,
  M.~Andreetto, and H.~Adam,
\newblock ``Mobilenets: Efficient convolutional neural networks for mobile
  vision applications,''
\newblock {\em CoRR}, vol. abs/1704.04861, 2017.


\bibitem{sahbi00004} H. Sahbi. "Learning laplacians in chebyshev graph convolutional networks." Proceedings of the IEEE/CVF International Conference on Computer Vision. 2021.




\bibitem{DBLP:conf/icml/TanL19}
M.~Tan and Q.~V. Le,
\newblock ``Efficientnet: Rethinking model scaling for convolutional neural
  networks,''
\newblock in {\em {ICML}}. 2019, vol.~97, {PMLR}.


\bibitem{DBLP:journals/corr/HintonVD15}
G.~E. Hinton, O.~Vinyals, and J.~Dean,
\newblock ``Distilling the knowledge in a neural network,''
\newblock {\em CoRR}, vol. abs/1503.02531, 2015.



\bibitem{sahbi00006} A. Mazari and H. Sahbi. "MLGCN: Multi-Laplacian graph convolutional networks for human action recognition." The British Machine Vision Conference (BMVC). 2019.



\bibitem{DBLP:conf/iclr/ZagoruykoK17}
S.~Zagoruyko and N.~Komodakis,
\newblock ``Paying more attention to attention: Improving the performance of
  convolutional neural networks via attention transfer,''
\newblock in {\em {ICLR}}, 2017.

\bibitem{DBLP:journals/corr/RomeroBKCGB14}
A.~Romero, N.~Ballas, S.~Ebrahimi Kahou, A.~Chassang, C.~Gatta, and Y.~Bengio,
\newblock ``Fitnets: Hints for thin deep nets,''
\newblock in {\em {ICLR}}, 2015.

\bibitem{DBLP:conf/aaai/MirzadehFLLMG20}
S.{-}I. Mirzadeh, M.~Farajtabar, A.~Li, N.~Levine, A.~Matsukawa, and
  H.~Ghasemzadeh,
\newblock ``Improved knowledge distillation via teacher assistant,''
\newblock in {\em {AAAI}}, 2020.



\bibitem{Sahbi2015} H.~Sahbi, ``Imageclef annotation with explicit context-aware kernel maps,''   \emph{International Journal of Multimedia Information Retrieval}, pp.  113--128, 2015.



\bibitem{DBLP:conf/cvpr/ZhangXHL18}
Y.~Zhang, T.~Xiang, T.~M. Hospedales, and H.~Lu,
\newblock ``Deep mutual learning,''
\newblock in {\em {CVPR}}, 2018.

\bibitem{DBLP:conf/cvpr/AhnHDLD19}
S.~Ahn, S.~X. Hu, A.~C. Damianou, N.~D. Lawrence, and Z.~Dai,
\newblock ``Variational information distillation for knowledge transfer,''
\newblock in {\em {CVPR}}, 2019.



\bibitem{ref00004555} S. Tollari,  P. Mulhem, M. Ferecatu, H. Glotin, M. Detyniecki, P. Gallinari, H. Sahbi and Z-Q. Zhao. "A comparative study of diversity methods for hybrid text and image retrieval approaches." In Workshop of the Cross-Language Evaluation Forum for European Languages, pp. 585-592. Springer, Berlin, Heidelberg, 2008.


\bibitem{DBLP:conf/nips/CunDS89}
Y.~LeCun, J.~S. Denker, and S.~A. Solla,
\newblock ``Optimal brain damage,''
\newblock in {\em {NIPS}}, 1989.

\bibitem{DBLP:conf/nips/HassibiS92}
B.~Hassibi and D.~G. Stork,
\newblock ``Second order derivatives for network pruning: Optimal brain
  surgeon,''
\newblock in {\em {NIPS}}, 1992.


\bibitem{DBLP:conf/nips/HanPTD15}
S.~Han, J.~Pool, J.~Tran, and W.~J. Dally,
\newblock ``Learning both weights and connections for efficient neural
  network,''
\newblock in {\em {NIPS}}, 2015.


\bibitem{sahbiicpr2002} H. Sahbi and N. Boujemaa. "Coarse-to-fine support vector classifiers for face detection." Object recognition supported by user interaction for service robots. Vol. 3. IEEE, 2002.


\bibitem{DBLP:journals/corr/HanMD15}
S.~Han, H.~Mao, and W.~J. Dally,
\newblock ``Deep compression: Compressing deep neural network with pruning,
  trained quantization and huffman coding,''
\newblock in {\em {ICLR}}, 2016.

\bibitem{DBLP:conf/iclr/0022KDSG17}
H.~Li, A.~Kadav, I.~Durdanovic, H.~Samet, and H.~P. Graf,
\newblock ``Pruning filters for efficient convnets,''
\newblock in {\em {ICLR}}, 2017.


\bibitem{sahbithesis} H. Sahbi,  Coarse-to-fine support vector machines for hierarchical face detection. Diss. PhD thesis, Versailles University, 2003.



\bibitem{DBLP:conf/iccv/LiuLSHYZ17}
Z.~Liu, J.~Li, Z.~Shen, G.~Huang, S.~Yan, and C.~Zhang,
\newblock ``Learning efficient convolutional networks through network
  slim,''
\newblock {\em {ICCV}}. 2017.

\bibitem{howard2019}
Howard, Andrew, et al. "Searching for mobilenetv3." ICCV 2019.

\bibitem{LiLi2020}
Li, Yingwei, et al. "Neural architecture search for lightweight non-local networks." In CVPR 2020.

\bibitem{LeLe2022} Le, Huu, et al. "AdaSTE: An Adaptive Straight-Through Estimator to Train Binary Neural Networks." In CVPR 2022.

\bibitem{sahbijmlr06} H. Sahbi, D. Geman. A hierarchy of support vector machines for pattern detection. Journal of Machine Learning Research 7.Oct (2006): 2087-2123.




\bibitem{refref74} B. Koneru, N. Girish, and V. Vasudevan. ”Sparse artificial neural networks using a novel smoothed LASSO penalization.”IEEE TCS II: Express Briefs 66.5 (2019): 848-852.

\bibitem{Tu2018} J.  Tu, M.  Liu, and H.  Liu, “Skeleton-based human action recognition using spatial temporal 3d convolutional neural networks,” in Proc. IEEE Int. Conf. Multimedia Expo (ICME), San Diego, CA, United states, Jul. 2018, pp. 1–6.


\bibitem{sahbi00008} H. Sahbi and N. Boujemaa. "From coarse to fine skin and face detection." Proceedings of the eighth ACM international conference on Multimedia. 2000.




\bibitem{ref00001244}  T. Stefan, H. Sahbi and M. Steinebach. "Using entropy for image and video authentication watermarks." Security, Steganography, and Watermarking of Multimedia Contents VIII. Vol. 6072. SPIE, 2006.

\bibitem{LiuLiu2018} H.  Liu,  J. Tu,  M.  Liu,  and R.  Ding,  “Learning explicit shape and motion evolution maps for skeleton-based human action recognition,” in Proc.  IEEE Int.  Conf.  Acoust.  Speech Signal Process. (ICASSP), Calgary,  AB,  Canada,  Apr.  2018,   pp. 1333–1337.

\bibitem{ref000010}  H. Sahbi. "CNRS-TELECOM ParisTech at ImageCLEF 2013 Scalable Concept Image Annotation Task: Winning Annotations with Context Dependent SVMs." CLEF (Working Notes). 2013.

 \bibitem{Ji2014}
Y. Ji, G. Ye, and H. Cheng. Interactive body part contrast mining for human interaction recognition. In IEEE International Conference on Multimedia and Expo Workshop, pages 1–6. IEEE, 2014.

\bibitem{Li2015a}
W. Li, L. Wen, M. Choo Chuah, and S. Lyu. Category-blind human action recognition: A practical recognition system. In IEEE International Conference on Computer Vision, pages 4444–4452, 2015.



\bibitem{ref000012} H. Sahbi, J-Y. Audibert, R. Keriven. "Graph-cut transducers for relevance feedback in content based image retrieval." 2007 IEEE 11th International Conference on Computer Vision. IEEE, 2007.

 
\bibitem{Yun2012}
K.  Yun, J.  Honorio, D.  Chattopadhyay,  T.L.  Berg, and D. Samaras. Two-person interaction detection using body pose features and multiple instance learning. In IEEE Conference on Computer Vision and Pattern Recognition Workshop, 2012.

 
\bibitem{Liu2021}
Liu, Jianbo, Ying Wang, Shiming Xiang, and Chunhong Pan. "Han: An efficient hierarchical self-attention network for skeleton-based gesture recognition." arXiv preprint arXiv:2106.13391 (2021).



\bibitem{ref000014}   T. Napoléon and H. Sahbi. "From 2D silhouettes to 3D object retrieval: contributions and benchmarking." EURASIP Journal on Image and Video Processing 2010 (2010): 1-17.



\bibitem{RiemannianManifoldTraject}
A. Kacem, M. Daoudi, B. Ben Amor, S. Berretti, J-Carlos.  Alvarez-Paiva. A Novel Geometric Framework on Gram Matrix Trajectories for Human Behavior Understanding.  IEEE Transactions on Pattern Analysis and Machine Intelligence, 28 September 2018


\bibitem{Nguyen2019}
XS. Nguyen, L. Brun, O. Lezoray and S. Bougleux. A neural network based on SPD manifold learning for skeleton-based hand gesture recognition. In IEEE International Conference on Computer Vision and Pattern Recognition, CVPR, 2019.

\bibitem{ref20}
Z. Huang, C. Wan, T. Probst, and L. V. Gool. Deep Learning on Lie Groups for Skeleton-Based Action Recognition. In CVPR, pages 6099–6108, 2017.


\bibitem{ref000014a}  M. Jiu and H. Sahbi. "Context-aware deep kernel networks for image annotation." Neurocomputing 474 (2022): 154-167.

\bibitem{Shi2018}
L. Shi, Y. Zhang, J. Cheng, and H. Lu. Non-Local Graph Convolutional Networks for Skeleton-Based Action Recognition. CoRR, abs/1805.07694, 2018

\bibitem{Li2019}
B. Li, X. Li, Z. Zhang, and F. Wu. Spatio-temporal graph routing for skeleton-based action recognition. AAAI Conference on Artificial Intelligence, 2019


\bibitem{ref000014b}   H. Sahbi.  "Topologically-Consistent Magnitude Pruning for Very Lightweight Graph Convolutional Networks." In 2022 IEEE International Conference on Image Processing (ICIP), pp. 3495-3499. IEEE, 2022.

\bibitem{Jiang2020}
T. Jiang, T. Huang and Y. Tian. Global Co-occurrence Feature Learning and Active Coordinate System Conversion for Skeleton-based Action Recognition. In Winter Conference on Applications of Computer Vision, pages 586-594. IEEE, 2020.

\bibitem{Yanc2018}
 S. Yan, Y. Xiong,  and D.  Lin.   Spatial Temporal Graph Convolutional Networks for Skeleton-Based Action Recognition. In AAAI,  pages 7444-7452, 2018.


\bibitem{Wen2019}
Y. Wen, L. Gao, H. Fu, F. Zhang, and S. Xia. Graph CNNs with motif and variable temporal block for skeleton-based action recognition. AAAI Conference on Artificial Intelligence, 2019.



\bibitem{ref000014c}  H.  Sahbi and H. Zhan. "FFNB: Forgetting-Free Neural Blocks for Deep Continual Learning." In The British Machine Vision Conference (BMVC). 2021.

\bibitem{Liu2018}
M. Liu and J. Yuan. Recognizing human actions as the evolution of pose estimation maps. In IEEE Conference on Computer Vision and Pattern Recognition, 2018.


\bibitem{ref32} M.  Liu,  H.  Liu,  and C.  Chen. Enhanced Skeleton Visualization for View Invariant Human Action Recognition. Pattern Recognition, 68:346–362, 2017.



\bibitem{ref000014d}  M.  Jiu and H. Sahbi. "DHCN: Deep hierarchical context networks for image annotation." In ICASSP 2021-2021 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP), pp. 3810-3814. IEEE, 2021.


\bibitem{Lee2017}
I. Lee, D. Kim, S. Kang, and S. Lee. Ensemble deep learning for skeleton-based action recognition using temporal sliding lstm networks. In IEEE International Conference on Computer Vision, pages 1012–1020, 2017.


\bibitem{Zhang2017b}
S. Zhang, X. Liu, and J. Xiao. On geometric features for skeleton-based action recognition using multilayer lstm networks. In Winter Conference on Applications of Computer Vision, pages 148–157. IEEE, 2017

\bibitem{Zhang2017}
P. Zhang, C. Lan, J. Xing, W. Zeng, J. Xue, and N. Zheng. View adaptive recurrent neural networks for high performance human action recognition from skeleton data. arXiv preprint arXiv:1703.08274, 2017.


\bibitem{ref000014e} L. Wang and H. Sahbi. "Directed acyclic graph kernels for action recognition." In Proceedings of the IEEE International Conference on Computer Vision, pp. 3168-3175. 2013.
  

\bibitem{ref44}
A. Shahroudy, J. Liu, T. T. Ng, and G. Wang. NTU RGB+D: A Large Scale Dataset for 3D Human Activity Analysis. In CVPR, pages 1010–1019, 2016.

  \bibitem{ref31}
J. Liu, G. Wang, P. Hu, L.-Y. Duan, and A. C. Kot. Global Context-Aware Attention LSTM Networks for 3D Action Recognition. In CVPR, pages 3671–3680, 2017.


\bibitem{ref000014f}  M. Jiu and H.  Sahbi. "Deep representation design from deep kernel networks." Pattern Recognition 88 (2019): 447-457.



  \bibitem{ref51} H. Wang and L. Wang. Modeling Temporal Dynamics and Spatial Configurations of Actions Using Two-Stream Recurrent Neural Networks. CVPR, pages 3633–3642, 2017.


 \bibitem{ref36}
  J. C. Nez, R. Cabido, J. J. Pantrigo, A. S. Montemayor, and J. F. Vlez. Convolutional Neural Networks and Long Short-Term Memory for Skeleton-based Human Activity and Hand Gesture Recognition. Pattern Recognition, 76(C):80–94, 2018.



\bibitem{ref000014g}  H.  Sahbi.  "Deep total variation support vector networks." In Proceedings of the IEEE/CVF International Conference on Computer Vision Workshops, pp. 0-0. 2019.


\bibitem{ref53}
 P. Wang, Z. Li, Y. Hou, and W. Li. Action Recognition Based on Joint Trajectory Maps Using Convolutional Neural Networks. In ACM MM, pages 102–106, 2016.


\bibitem{ke2017}
Q. Ke, M. Bennamoun, S. An, F. Sohel, and F. Boussaid. A new representation of skeleton sequences for 3D action recognition. In IEEE Conference on Computer Vision and Pattern Recognition Workshop, pages 4570–4579. IEEE, 2017.



\bibitem{ref000014h}  Q. Oliveau and H. Sahbi. "From transductive to inductive semi-supervised attributes for ship category recognition." In IGARSS 2018-2018 IEEE International Geoscience and Remote Sensing Symposium, pp. 4827-4830. IEEE, 2018.


 \bibitem{REF1}  X. Chen, H. Guo, G. Wang, and L. Zhang, “Motion feature augmented recurrent neural network for skeleton-based dynamic hand gesture recognition,” in Proc. Int. Conf. Image Process. (ICIP), Beijing, China, Sep. 2017, pp. 2881–2885.



\bibitem{DeepGRU}
M. Maghoumi, JJ. LaViola Jr. DeepGRU: Deep Gesture Recognition Utility. In arXiv preprint arXiv:1810.12514, 2018



\bibitem{ref000014i}   Q.  Oliveau and H. Sahbi. "Semi-supervised deep attribute networks for fine-grained ship category recognition." In IGARSS 2018-2018 IEEE International Geoscience and Remote Sensing Symposium, pp. 6871-6874. IEEE, 2018.

\bibitem{GCALSTM}
J. Liu, G. Wang, L. Duan, K. Abdiyeva, and A. C. Kot. Skeleton-based human action recognition with global context-aware attention lstm networks. IEEE Transactions on Image Processing, 27(4):1586–1599, April 2018


\bibitem{Liu2016}
J. Liu, A. Shahroudy, D. Xu, and G. Wang. Spatio-temporal LSTM with trust gates for 3D human action recognition. In European Conference on Computer Vision, pages 816–833. Springer, 2016




\bibitem{ref000014j}  H.  Sahbi.  "Learning cca representations for misaligned data." In Proceedings of the European Conference on Computer Vision (ECCV) Workshops, pp. 0-0. 2018.


\bibitem{Wangb2012}
  J. Wang, Z. Liu, Y. Wu, and J. Yuan. Mining Actionlet Ensemble for Action Recognition with Depth Cameras. In CVPR, pages 1290–1297, 2012.


\bibitem{Yang2014}
X. Yang and Y. Tian. Effective 3D action recognition using eigenjoints. Journal of Visual Communication and Image Representation, 25(1):2–11, 2014.





\bibitem{ref000014k}  Q.  Oliveau and H.  Sahbi. "Semantic-free attributes for image classification." In 2016 23rd International Conference on Pattern Recognition (ICPR), pp. 1577-1582. IEEE, 2016.

\end{thebibliography}

\documentclass[12pt,journal,compsoc,onecolumn]{IEEEtran}
\usepackage{amsmath,graphicx,cleveref,caption,subcaption,dsfont,multirow,multicol,amssymb,amsfonts,mathtools,pifont}

 \usepackage{color}

\usepackage{booktabs}
\usepackage{lipsum}
\usepackage{bm}
\usepackage{stmaryrd}
\usepackage{algpseudocode}
\usepackage[ruled,vlined]{algorithm2e}
\usepackage[noadjust]{cite}
\let\OLDthebibliography\thebibliography
\renewcommand\thebibliography[1]{
  \OLDthebibliography{#1}
  \setlength{\parskip}{0pt}
  \setlength{\itemsep}{1.47pt plus 0.3ex}
}

\newcommand{\cmark}{\ding{51}}
\newcommand{\xmark}{\ding{55}}

\def\UU{{\bf U}}

\def\U{{\mathbf U}} 
\def\V{{\mathbf V}} 
\def\A{{\mathbf \Lambda}}

\def\A{{\bf A}}

\def\U{{\bf U}}
\def\W{{\bf W}}
\def\WW{{\bf W}}
\def\SS{{\cal S}}

\def\N{{\cal N}}
\def\G{{\cal G}}
\def\V{{\cal V}}
\def\E{{\cal E}}
\def\F{{\cal F}}

\def\M{{\bf M}}

\title{Miniaturized Graph Convolutional Networks with Topologically Consistent Pruning}
\author{Hichem Sahbi \\ 
Sorbonne University, UPMC, CNRS, LIP6, France}

\begin{document}
\maketitle
\begin{abstract}
  Magnitude pruning  is one of the mainstream  methods in lightweight architecture design whose goal is to extract subnetworks  with the largest weight connections. This method is known to be successful, but under very high pruning regimes, it suffers from topological inconsistency which renders the extracted subnetworks disconnected, and this hinders their generalization ability. In this paper, we devise a novel magnitude pruning method that allows extracting subnetworks while guarantying their topological consistency. The latter ensures that only accessible and co-accessible --- impactful --- connections  are kept in the resulting lightweight networks. Our solution is based on a novel reparametrization and two supervisory bi-directional networks which implement accessibility/co-accessibility and guarantee that only connected subnetworks will be selected during training. This solution allows enhancing generalization significantly, under very high pruning regimes, as corroborated through extensive experiments, involving graph convolutional networks, on the challenging task of skeleton-based action recognition.\\
  
 {\noindent {\bf Keywords.}  Graph convolutional networks, lightweight design, magnitude pruning, skeleton-based recognition}
\end{abstract}

\section{Introduction}

Deep convolutional networks are nowadays becoming mainstream in solving many pattern classification tasks  including image and action recognition \cite{Krizhevsky2012,refref10}. Their principle consists in training convolutional filters together with pooling and attention  mechanisms that maximize classification performances.   Many existing  convolutional networks were initially dedicated to grid-like data, including images \cite{He2016,He2017,HuangCVPR2017,Ronneberger2015}.  However,  data sitting on top of irregular domains (such as skeleton graphs  in action recognition) require extending    convolutional networks  to  general graph structures, and these extensions are known as graph convolutional networks (GCNs) \cite{Zhua2016,zhang2020}. Two families of GCNs exist in the literature: spectral and spatial. Spectral methods are based on graph Fourier transform  \cite{spectral1997,kipf17,Levie2018,Li2018,Bresson16,Bruna2013,Henaff2015,sahbi00006} while spatial ones rely on message passing and attention \cite{Gori2005,Micheli2009,Wu2019,Hamilton2017}. Whilst spatial GCNs have been relatively more effective compared to spectral ones, their precision is reliant on the attention matrices that capture context and node-to-node relationships \cite{attention2019}.  With multi-head attention, GCNs are more accurate but overparametrized and computationally overwhelming.  \\
\indent Many solutions are proposed in the literature to reduce  time and  memory  footprint of  convolutional networks including GCNs~\cite{DBLP:conf/cvpr/HuangLMW18,DBLP:conf/cvpr/SandlerHZZC18,DBLP:journals/corr/HowardZCKWWAA17,DBLP:conf/icml/TanL19}. Some of them pretrain oversized networks prior to reduce their computational complexity  (using distillation \cite{DBLP:journals/corr/HintonVD15,DBLP:conf/iclr/ZagoruykoK17,DBLP:journals/corr/RomeroBKCGB14,DBLP:conf/aaai/MirzadehFLLMG20,DBLP:conf/cvpr/ZhangXHL18,DBLP:conf/cvpr/AhnHDLD19}, linear algebra \cite{howard2019},  quantization \cite{DBLP:journals/corr/HanMD15} and pruning~\cite{DBLP:conf/nips/CunDS89,DBLP:conf/nips/HassibiS92,DBLP:conf/nips/HanPTD15}), whilst others build efficient networks from scratch using neural architecture search~\cite{LiLi2020}.  In particular, pruning methods, either unstructured or structured are  currently mainstream, and their principle consists in removing connections whose impact on the classification performance is the least noticeable.  Unstructured pruning \cite{DBLP:conf/nips/HanPTD15,DBLP:journals/corr/HanMD15} proceed by dropping out connections individually using different proxy criteria,  such as weight magnitude,  and then retrain the resulting pruned networks.   In contrast,  structured pruning \cite{DBLP:conf/iclr/0022KDSG17,DBLP:conf/iccv/LiuLSHYZ17} removes groups of connections,  entire filters or subnetworks using different  mechanisms such as grouped sparsity.   However,  existing pruning methods either structured or unstructured suffer from several drawbacks.  On the one hand, structured pruning may reach high speedup on usual hardware, but its downside resides in the rigidity of the class of learnable networks.  On the other hand, unstructured pruning is more flexible, but its discrimination is limited at high pruning regimes due to {\it topological disconnections}, and handling the latter is highly intractable as adding or removing any connection {\it combinatorially} affects the others. \\
\indent As contemporary network sizes grow into billions of parameters, studying high compression regimes has been increasingly important on very large networks. Nevertheless, pruning relatively smaller networks is even more challenging as this usually leads to highly disconnected and untrainable subnetworks, even at reasonably (not very) large pruning rates. Hence, we target our contribution towards mid-size  networks including GCNs in order to fit not only the usual edge devices, such as smartphones, but also highly {\it miniaturized} devices endowed with very limited computational resources (e.g., smart glasses). Considering the aforementioned issues, we introduce in this paper a new lightweight network design which guarantees the topological consistency of the extracted subnetworks.  Our proposed solution is variational and proceeds by training pruning masks and weight parameters that maximize classification performances while guaranteeing the {\it accessibility} of the unpruned connections  (i.e., their reachability from the network input) and their {\it co-accessibility} (i.e., their actual contribution in the evaluation of the output). Hence, only topologically consistent subnetworks are considered when selecting connections. Extensive experiments, on the challenging task of skeleton-based action recognition, show the outperformance of our proposed topologically consistent pruning.
\def\SS{{\cal S}}
\def\G{{\cal G}}
\def\V{{\cal V}}
\def\E{{\cal E}}
\def\A{{\bf A}}
\def\F{{\cal F}}
\def\W{{\bf W}}
\def\U{{\bf {W}}}
\def\UU{{\bf U}} 
\def\N{{\cal N}}
\def\WW{{\bf W}}
\def\M{{\bf M}}
\def\SSS{{\bf S}}

\section{A Glimpse on GCNs}
Let $\SS=\{\G_i=(\V_i, \E_i)\}_i$ denote a collection of graphs with $\V_i$, $\E_i$ being respectively the nodes and the edges of $\G_i$. Each graph $\G_i$ (denoted for short as $\G=(\V, \E)$) is endowed with a signal $\{\psi(u) \in \mathbb{R}^s: \ u \in \V\}$ and associated with an adjacency matrix $\A$ with each entry  $\A_{uu'}>0$ iff $(u,u') \in \E$ and $0$ otherwise. GCNs aim at learning a set of $C$ filters $\F$ that define convolution on $n$ nodes of $\G$ (with $n=|\V|$) as $(\G \star \F)_\V = f\big(\A \  \UU^\top  \   \W\big)$, here $^\top$ stands for transpose,  $\UU \in \mathbb{R}^{s\times n}$  is the  graph signal, $\W \in \mathbb{R}^{s \times C}$  is the matrix of convolutional parameters corresponding to the $C$ filters and  $f(.)$ is a nonlinear activation applied entrywise. In   $(\G \star \F)_\V$, the input signal $\UU$ is projected using $\A$ and this provides for each node $u$, the  aggregate set of its neighbors. Entries of $\A$ could be handcrafted or learned so  $(\G \star \F)_\V$ implements a convolutional block with two layers; the first one aggregates signals in $\N(\V)$ (sets of node neighbors) by multiplying $\UU$ with $\A$ while the second layer achieves convolution by multiplying the resulting aggregates with the $C$ filters in $\W$. Learning  multiple adjacency (also referred to as attention) matrices (denoted as $\{\A^k\}_{k=1}^K$) allows us to capture different contexts and graph topologies when achieving aggregation and convolution.  With multiple matrices $\{\A^k\}_k$ (and associated convolutional filter parameters $\{\W^k\}_k$),   $(\G \star \F)_\V$ is updated as  $f\big(\sum_{k=1}^K \A^k   \UU^\top     \W^k\big)$. Stacking aggregation and convolutional layers, with multiple  matrices $\{\A^k\}_k$, makes GCNs accurate but heavy. We propose  subsequently a method that makes our networks lightweight and still effective.

 \section{Relaxed Magnitude Pruning}

 In the rest of this paper,   a given GCN is  subsumed as a multi-layered neural network $g_\theta$  whose weights defined as $\theta =  \left\{\WW^1,\dots, \WW^L \right\}$, with $L$ being its depth,  $\WW^\ell \in \mathbb{R}^{d_{\ell-1} \times d_{\ell}}$ its $\ell^\textrm{th}$  layer weight tensor, and $d_\ell$ the dimension of $\ell$. The output of a given layer  $\ell$ is defined as $\mathbf{\phi}^{\ell} = f_\ell({\WW^\ell}^\top \  \mathbf{\phi}^{\ell-1})$, $\ell \in \{2,\dots,L\}$, being $f_\ell$ an activation function. Without a loss of generality, we omit the bias in the definition of $\mathbf{\phi}^{\ell}$. Magnitude Pruning (MP) consists in zeroing the smallest weights in $g_\theta$ (up to a pruning rate),  while retraining the remaining weights. A relaxed variant of MP is obtained by multiplying $\WW^\ell$ with a differentiable mask $\psi(\WW^\ell)$ applied entrywise to $\WW^\ell$. The entries of  $\psi(\WW^\ell)$ are set depending on whether the underlying layer connections are kept or removed, so $\mathbf{\phi}^{\ell} = f_\ell( (\WW^\ell \odot \psi(\WW^\ell))^\top \ \mathbf{\phi}^{\ell-1} )$, here $\odot$ stands for the element-wise matrix product. In this definition, $\psi({\WW}^\ell)$ enforces the prior that smallest weights should be removed from the network. In order to achieve magnitude pruning,  $\psi$ must be symmetric,  bounded in $[0,1]$, and $\psi(\omega) \leadsto 1$ when $|\omega|$ is sufficiently large and $\psi(\omega) \leadsto 0$ otherwise\footnote{A possible choice, used in practice, that satisfies these four conditions is $\psi(\omega)=2\sigma(\omega^2)-1$ with $\sigma$ being the sigmoid function.}.\\

\indent Pruning is achieved using a global loss as a combination of a cross entropy term denoted as ${\cal L}_e$, and a budget cost which  measures the difference between the targeted cost (denoted as $c$) and the actual number of unpruned connections
\begin{equation}\label{eq3}
 \displaystyle \min_{\{{\WW}^\ell\}_\ell}  \displaystyle {\cal L}_e\big(\{{\WW}^\ell \odot \psi({\WW}^\ell)\}_\ell\big) +  \lambda \big(\sum_{\ell=1}^{L-1}  {\bf 1}_{d_\ell}^\top \psi({\WW}^\ell) {\bf 1}_{d_{\ell+1}}  - c\big)^2, 
    \end{equation}
\noindent here ${\bf 1}_{d_\ell}$ denotes a vector of $d_\ell$ ones. When $\lambda$ is sufficiently large, Eq.~\ref{eq3} focuses on minimizing the budget loss while progressively making $\{\psi({\WW}^\ell)\}_\ell$ crisp (almost binary) using annealing. As training evolves, the right-hand side term reaches its minimum and stabilizes while the gradient of the global loss becomes dominated by the gradient of the left-hand side term, and this maximizes further the classification performances. 

\section{Topologically Consistent Magnitude Pruning} 
The aforementioned pruning formulation is relatively effective (as shown later in experiments), however, it suffers from several drawbacks. On the one hand, removing connections independently may result into {\it topologically inconsistent} networks (see section~\ref{section3}), i.e., either completely disconnected or having isolated connections. On the other hand, high pruning rates may lead to an over-regularization effect and hence weakly discriminant lightweight networks, especially when the latter include isolated connections (see again later experiments). In what follows, we introduce a more principled pruning framework that guarantees the topological consistency of the pruned networks and allows improving generalization even at very high pruning rates. 

\subsection{Accessibility and Co-accessibility}\label{section3}

Our formal definition of topological consistency relies on two principles: {\it accessibility and co-accessibility} of connections in $g_\theta$ \cite{ref000014b}. Let's remind $\psi(\U_{ij}^\ell)$ as crisped (binary) function that indicates the presence or absence of a connection between the i-th and the j-th neurons of layer $\ell$. This  connection is referred to as accessible if $ \ \exists i_1,\dots,i_{\ell-1}$, s.t. $\psi(\U_{i_1,i_2}^1)=\dots =\psi(\U_{i_{\ell-1},i}^{\ell-1})=1$, and it is co-accessible if $\exists i_{\ell+1},\dots,i_L$, s.t. $\psi(\U_{j,i_{\ell+1}}^{\ell+1})= \dots =\psi(\U_{i_{L-1},i_L}^{L})=1$.  

Considering $\SSS_a^\ell = \psi(\U^1) \ \psi(\U^2) \dots \psi(\U^{\ell-1})$ and $\SSS_c^\ell = \psi(\U^{\ell+1}) \ \psi(\U^{\ell+2}) \dots \psi(\U^{L})$, and following the above definition, it is easy to see that a connection between $i$ and $j$ is accessible (resp. co-accessible) iff the i-th column (resp. j-th row) of $\SSS_a^\ell$ (resp. $\SSS_c^\ell$) is different from the null vector. A network is called {topologically consistent} iff all its connections are both accessible and co-accessible. Accessibility guarantees that incoming connections to the i-th neuron carry out effective activations resulting from the evaluation of $g_\theta$ up to layer $\ell$. Co-accessibility is equivalently important and guarantees that the outgoing activation from the j-th neuron actually contributes in the evaluation of the network output. A connection --- not satisfying accessibility or co-accessibility and even when its magnitude is large --- becomes useless and should be removed when $g_\theta$ is pruned.\\  \indent For any given network, parsing all its topologically consistent subnetworks and keeping only the one that minimizes Eq.~\ref{eq3} is highly combinatorial. Indeed, the accessibility of a given connection depends on whether its  preceding and subsequent ones are kept or removed, and any masked connections may affect the accessibility of the others. In what follows, we introduce a solution that prunes a given network while guaranteeing its topological consistency. 

\subsection{(Co)Accessibility Networks and Loss Function}
Our solution relies on two supervisory networks that measure  accessibility and co-accessibility of connections in $g_\theta$. These two networks, denoted as $\phi_r$  and $\phi_l$, have exactly the same architecture as $g_\theta$ with only a few differences: indeed, $\phi_r$ measures accessibility and inherits the same connections in $g_\theta$ with the only difference that their weights correspond to  $\{\psi(\U^\ell)\}_\ell$ instead of $\{\U^\ell \odot \psi(\U^\ell)\}_\ell$. Similarly, $\phi_l$ inherits the same connections and weights as $\phi_r$, however these connections are reversed in order to measure accessibility in the opposite direction (i.e., co-accessibility). Note that weights $\{\U^\ell\}_\ell$ are shared across all the networks $g_\theta$, $\phi_r$ and $\phi_l$.\\
\noindent Considering the definition of accessibility and co-accessibility, one may define layerwise outputs $\phi_r^{\ell} := h\big(\psi({\U_{\ell-1}})^\top \ \phi_r^{\ell-1}\big)$, and $\phi_l^{\ell} :=  h\big(\psi({\U_{\ell}})  \ \phi_l^{\ell+1}\big)$,  being  ${\phi_r^1} = {\bf 1}_{d_1}$, ${\phi_l^L} = {\bf 1}_{d_L}$, ${\bf 1}_{d_1}$ the vector of $d_1$ ones and $h$ the Heaviside activation. With  $\phi_r^{\ell}$ and $\phi_l^{\ell}$, non-zero entries of the matrix $(\phi^{\ell} _r \phi^{{\ell+1}^\top}_l) \odot \psi({\WW}^\ell) $ correspond to selected connections in $g_\theta$ which are also accessible and co-accessible. By plugging  this matrix into Eq.~\ref{eq3}, we redefine our topologically consistent pruning loss 
\begin{equation}\label{eq4}
  \begin{array}{ll}
  \displaystyle \min_{\{{\WW}^\ell\}_\ell} & \displaystyle {\cal L}_e\big(\{{\WW}^\ell \odot \psi({\WW}^\ell) \odot \phi^{\ell}_r \phi^{{\ell+1}^\top} _l \}_\ell\big) \ + \  \lambda \big(\sum_{\ell=1}^{L-1}  \phi^{{\ell}^\top} _r  \psi({\WW}^\ell) \phi^{\ell+1}_l - c\big)^2,
\end{array}
     \end{equation}
     \begin{equation}\label{eq7}
  \begin{array}{lllll}
 \textrm{with} &    \phi_r^{\ell} &:=& h\big((\phi^{\ell-1}_r \phi^{{\ell}^\top} _l \odot \psi({\U_{\ell-1}}))^\top  \ \phi_r^{\ell-1}\big)  & \\
 &     & &  &    \\ 
    &\phi_l^{\ell} &:=&  h\big((\phi^{\ell}_r \phi^{{\ell+1}^\top} _l \odot \psi({\U_{\ell}})) \phi^{\ell+1}_l\big).   &  \\
\end{array}
\end{equation}
It is clear that accessibility networks in Eq.~\ref{eq7} cannot be modeled using standard feedforward neural networks, so more complex (highly recursive and interdependent) networks should be considered which also leads to exploding gradient.  In order to make Eq.~\ref{eq7} simpler and still trainable with standard feedforward networks, we constrain entries of $\psi({\U_{\ell}})$ to take non-zero values iff the underlying connections are kept and accessible/co-accessible; in other words, $\phi^{{\ell}^\top} _r \psi({\U_{\ell}}) \phi^{\ell+1}_l$ should approximate $ {\bf 1}_{d_\ell}^{\top} \psi({\U_{\ell}}){\bf 1}_{d_{\ell+1}}$ in order to guarantee that (i) unpruned connections are necessarily accessible/co-accessible and (ii) non accessible ones are necessarily pruned.  Hence, instead of~\ref{eq4} and~\ref{eq7}, a surrogate optimization problem is defined as
\begin{equation}\label{eq41}
  \begin{array}{ll}
  \displaystyle \min_{\{{\WW}^\ell\}_\ell} & \displaystyle {\cal L}_e\big(\{{\WW}^\ell \odot \psi({\WW}^\ell) \odot \phi^{\ell}_r \phi^{{\ell+1}^\top} _l \}_\ell\big) \ + \  \displaystyle \lambda \big(\sum_{\ell=1}^{L-1}  \phi^{\ell^\top} _r  \psi({\WW}^\ell) \phi^{\ell+1}_l - c\big)^2  \\
   & \ \ \ \ \  \ +  \displaystyle \eta \sum_{\ell=1}^{L-1}  \big[ {\bf 1}_{d_\ell}^\top \psi({\U_{\ell}}){\bf 1}_{d_{\ell+1}}   - {\phi^r_\ell}^\top  \psi(\U_\ell )  \phi_{\ell+1}^l \big],
\end{array}
     \end{equation}
 with now $\phi_r^{\ell} := h\big(\psi({\U_{\ell-1}})^\top \ \phi_r^{\ell-1}\big)$, $\phi_l^{\ell} :=  h\big(\psi({\U_{\ell}})  \ \phi_l^{\ell+1}\big)$. 

\subsection{Optimization}
Let ${\cal L}$ denote the global loss in Eq.~\ref{eq41}, the update of $\{{\WW}^\ell\}_\ell$ is achieved using stochastic gradient descent and by {\it simultaneously} backpropagating the gradients through the networks $g_\theta$, $\phi_r$ and $\phi_l$. More precisely, considering Eq.~\ref{eq41} and $\phi_r^{\ell}$, $\phi_l^{\ell}$, the gradient of the global loss w.r.t.   ${\WW}^\ell$ is obtained as
\begin{equation}\label{eq11}
  \frac{\partial {\cal L}}{\partial {\WW}^\ell}  + \sum_{k=\ell+1}^L \frac{\partial {\cal L}}{\partial \phi_r^{k}} \frac{\phi_r^{k}}{\phi_r^{k-1}} \dots  \frac{\partial \phi_r^{\ell+1}}{\partial {\WW}^\ell} + \sum_{k=1}^\ell \frac{\partial {\cal L}}{\partial \phi_l^{k}} \frac{\phi_l^{k}}{\phi_l^{k+1}} \dots  \frac{\partial \phi_l^{\ell}}{\partial {\WW}^\ell},
\end{equation} 
here the left-hand side term in Eq.~\ref{eq11} is obtained by backpropagating the gradient of $\cal L$ from the output to the input of the network $g_\theta$ whereas the mid terms are obtained by backpropagating the gradients of $\cal L$ from different layers to the input of $\phi_r$. In contrast, the right-hand side terms are obtained by backpropagating the gradients of $\cal L$ through $\phi_l$ in the opposite direction. Note that the evaluation of the gradients in  Eq.~\ref{eq11} relies on the straight through estimator (STE)~\cite{LeLe2022}; the sigmoid is used as a differentiable surrogate of $h$ during backpropagation while the initial Heaviside is kept when evaluating the responses of $\phi_r$, $\phi_l$ (i.e., forward steps). STE allows training differentiable accessibility networks while guaranteeing binary responses when evaluating these networks. 
\begin{table}[ht]
\centering \resizebox{0.9\columnwidth}{!}{
\begin{tabular}{ccccc}
{\bf Method} & {\bf Color} & {\bf Depth} & {\bf Pose} & {$ \ \ \ \ \  \ \ \ \ \ \  \  $ \bf Accuracy (\%) $ \ \ \ \  \ \ \ \ \ \ \  \  $}\\
\hline
  Two stream-color \cite{refref10}   & \cmark  &  \xmark  & \xmark  &  61.56 \\
Two stream-flow \cite{refref10}     & \cmark  &  \xmark  & \xmark  &  69.91 \\ 
Two stream-all \cite{refref10}      & \cmark  & \xmark   & \xmark  &  75.30 \\
\hline 
HOG2-depth \cite{refref39}        & \xmark  & \cmark   & \xmark  &  59.83 \\    
HOG2-depth+pose \cite{refref39}   & \xmark  & \cmark   & \cmark  &  66.78 \\ 
HON4D \cite{refref40}               & \xmark  & \cmark   & \xmark  &  70.61 \\ 
Novel View \cite{refref41}          & \xmark  & \cmark   & \xmark  &  69.21  \\ 
\hline
1-layer LSTM \cite{Zhua2016}        & \xmark  & \xmark   & \cmark  &  78.73 \\
2-layer LSTM \cite{Zhua2016}        & \xmark  & \xmark   & \cmark  &  80.14 \\ 
\hline 
Moving Pose \cite{refref59}         & \xmark  & \xmark   & \cmark  &  56.34 \\ 
Lie Group \cite{Vemulapalli2014}    & \xmark  & \xmark   & \cmark  &  82.69 \\ 
HBRNN \cite{Du2015}                & \xmark  & \xmark   & \cmark  &  77.40 \\ 
Gram Matrix \cite{refref61}         & \xmark  & \xmark   & \cmark  &  85.39 \\ 
TF    \cite{refref11}               & \xmark  & \xmark   & \cmark  &  80.69 \\  
\hline 
JOULE-color \cite{refref18}         & \cmark  & \xmark   & \xmark  &  66.78 \\ 
JOULE-depth \cite{refref18}         & \xmark  & \cmark   & \xmark  &  60.17 \\ 
JOULE-pose \cite{refref18}         & \xmark  & \xmark   & \cmark  &  74.60 \\ 
JOULE-all \cite{refref18}           & \cmark  & \cmark   & \cmark  &  78.78 \\ 
\hline 
Huang et al. \cite{Huangcc2017}     & \xmark  & \xmark   & \cmark  &  84.35 \\ 
Huang et al. \cite{ref23}           & \xmark  & \xmark   & \cmark  &  77.57 \\  
  \hline
Our  GCN baseline                  & \xmark  & \xmark   & \cmark  & \bf86.08                                                
\end{tabular}}
\caption{Comparison of our baseline GCN against related work on FPHA.}\label{compare2}
 \end{table} 
\begin{table}[ht]
 \begin{center}
\resizebox{0.9\columnwidth}{!}{
  \begin{tabular}{cccccc}
    \rotatebox{30}{Pruning rates}  & \rotatebox{30}{TC} &  \rotatebox{30}{\# parameters} &  \rotatebox{30}{\% of A-C}  &  \rotatebox{30}{Accuracy (\%)}  &  \rotatebox{30}{Observation}  \\
    \hline
    \hline
    0\%                   & NA & 1967616 & 100 &  86.08 &Baseline GCN \\ 
    \hline                                                  
   50.01\%  & \xmark  & 983549  & 98.74   & \bf85.04 &  MP w/o TC \\
        49.99\%                  &  \cmark  & 983836  & 100.0   & 84.34  &  MP w TC\\
        \hline                                                  
 75.21\%  & \xmark  & 487727   & 85.10   &\bf85.56 & MP w/o TC\\
           75.19\%                          &  \cmark  & 487990  & 100.0   & 85.21 & MP w TC \\
    \hline
    \hline
88.18\% & \xmark  & 232531   & 77.20   &  84.86 & MP w/o TC \\    
                 88.17\%                  &  \cmark  &  232768 & 100.0   & \bf85.91 & MP w TC \\
    \hline
  95.45\%                          & \xmark  & 89507   &64.34   & 83.47& MP w/o TC \\    
             95.45\%                   &  \cmark  & 89453  & 100.0   & \bf85.21 & MP w TC \\
    \hline
    98.99\%  & \xmark  & 19532    & 70.84 &80.00    & MP w/o TC\\
    99.01\%                               &  \cmark  & 19285  & 100.0   & \bf82.95 & MP w TC \\
    \hline
99.49\% & \xmark  & 7895  & 62.32   &  69.91&  MP w/o TC\\
     99.60\%                           &  \cmark  & 7732  & 100.0   & \bf80.00 & MP w TC
  \end{tabular}
}
\end{center}\caption{This table shows an ablation study, w and w/o Topological Consistency (TC), for different pruning rates on the FPHA dataset. We can see how Magnitude Pruning (MP), w/o TC, produces disconnected networks for high pruning rates, and this degrades performances, while MP with TC guarantees both Accessibility and Co-accessibility and also better generalization. \% A-C stands for percentage of Accessible and Co-accessible connections.}\label{table21}  
\end{table}

\section{Experiments}

We evaluate our GCNs on the task of action recognition using  the challenging First-Person Hand Action (FPHA) dataset~\cite{garcia2018}. This dataset consists of 1175 skeletons whose ground-truth includes 45 action categories with a high variability in style, speed and  scale as well as viewpoints. Each video, as a sequence of skeletons, is modeled with a graph $\G = (\V,\E)$ whose given node $v_j \in \V$ corresponds to the $j$-th hand-joint trajectory (denoted as $\{\hat{p}_j^t\}_t$)  and edge $(v_j, v_i) \in  \E$ exists iff the $j$-th and the $i$-th trajectories are spatially neighbors. Each trajectory in $\G$ is described using {\it temporal chunking} \cite{sahbiICPR2020}: this is obtained by first splitting the total duration of a video sequence into $M$ equally-sized temporal chunks ($M=32$ in practice), and assigning trajectory coordinates $\{\hat{p}_j^t\}_t$ to the $M$ chunks (depending on their time stamps), and then concatenating the averages of these chunks in order to produce the raw description (signal) of $v_j$.\\  

\noindent {\bf Implementation details and baseline GCN.} Our GCNs are trained end-to-end using Adam~\cite{Adam2014} for 2,700 epochs with a  momentum of $0.9$, batch size of $600$ and a global learning rate (denoted as $\nu(t)$) set depending on the change of the loss in Eq.~\ref{eq41}; when the latter increases (resp. decreases), $\nu(t)$  decreases as $\nu(t) \leftarrow \nu(t-1) \times 0.99$ (resp. increases as $\nu(t) \leftarrow \nu(t-1) \slash 0.99$). The mixing parameter $\eta$ in Eq.~\ref{eq41} is set to $1$ and $\lambda$ is slightly overestimated to $10$ in order to guarantee the implementation of the targeted pruning rates.  All these experiments are run on a GeForce GTX 1070 GPU (with 8 GB memory) and classification performances --- as average accuracy through action classes --- are evaluated using the protocol in~\cite{garcia2018} with 600 action sequences for training and 575 for testing. The architecture of our baseline GCN (taken from \cite{sahbiICPR2020}) consists of an attention layer of 16 heads applied to skeleton graphs whose nodes are encoded with 32-channels, followed by a convolutional layer of 128 filters, and a dense fully connected layer. This initial network is relatively heavy (for a GCN);  its includes 2 million parameters and it is accurate compared to the related work on the FPHA benchmark, as shown in Table~\ref{compare2}. Considering this GCN baseline, our goal is to make it lightweight while maintaining its high accuracy as much as possible.\\ 

\noindent {\bf Impact of TC on Lightweight CGNs.} We study the impact of  Topological Consistency (TC) on the performances of our lightweight GCNs for different pruning rates. Table.~\ref{table21} shows the positive impact of TC especially on highly pruned networks. This impact is less important (and sometimes negative) with low pruning regimes as the resulting networks have enough (a large number of) Accessible and Co-accessible (AC) connections, so having a few of these connections neither accessible nor co-accessible, i.e. removed, produces a well known regularization effect~\cite{dropconnect2013} that enhances performances. In contrast, with high pruning rates and without TC, this leads to over-regularized and very disconnected lightweight networks that suffer from under-fitting. With TC, both accessibility and co-accessibility  are guaranteed even with very high pruning regimes, and this also attenuates under-fitting, and ultimately improves generalization as again shown in table~\ref{table21}.

\section{Conclusion}
We introduce in this paper a novel lightweight network design based on magnitude pruning. The particularity of the method resides in its ability to select subnetworks with {\it only} accessible and co-accessible connections. The latter make the learned lightweight subnetworks topologically consistent and more accurate particularly at very high pruning regimes. The proposed approach relies on two supervisory networks, that implement accessibility and co-accessibility, which are trained simultaneously with the lightweight networks using a novel loss function. Extensive experiments, involving graph convolutional networks, on the challenging task of skeleton-based recognition show the substantial gain of our method.  Future work will investigate the  integration of this framework to different architectures and tasks involving extremely high pruning regimes.

\end{document}